OpenStudy (anonymous):

1 kg is equal to 103 hg. 103 g. 10-2 cg. 10-2 g.

OpenStudy (texaschic101):

1 kg = 10 hg 1 kg = 100 g 1 kg = 100,000 cg um...I am not sure @jim_thompson5910 .....whats up with this question ?

jimthompson5910 (jim_thompson5910):

1 kg = 1 kilogram = 1000 grams 1 hg = 1 hectogram = 100 grams 1 cg = 1 centigram = 1/100 of a gram (ie 100 cg = 1 g)

jimthompson5910 (jim_thompson5910):

oh i see what you're asking

OpenStudy (texaschic101):

I messed up on the 1 kg = 1000 g

jimthompson5910 (jim_thompson5910):

1 kg = 1000 grams 1 kg = 10^3 grams when the original poster copied/pasted the question, the exponent symbols didn't make it

jimthompson5910 (jim_thompson5910):

B should be 10^3 not 103

OpenStudy (texaschic101):

oh.....there exponents

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (texaschic101):

lol....I was so lost.......I am found now...thanks

jimthompson5910 (jim_thompson5910):

lol yw